Downton Ltd signs a non-cancellable five-year lease on an item of machinery on 1 July 2018. At the inception of the lease, the machinery has a fair value of $801,060.

The expected economic life of the machinery is six years, when it is expected to have a residual value of nil. There is a bargain purchase option that Downton Ltd, as the lessee, will be able to exercise at the end of the fifth year of the lease for $300,000.

There will be five annual payments of $200,000, the first being made on 30 June 2019. Included within each of the $200,000 payments is an amount of $25,000 which will be paid to the lessor to cover insurance and maintenance on the equipment. There were no initial direct costs associated with the lease. The equipment is to be depreciated on a straight-line basis.

The interest rate implicit in the lease is 12%.

Required:

  1. Prove that the interest rate implicit in the lease is 12%. Show workings.
  2. Prepare a schedule of lease payments for the duration of the lease in the following format.          Round to whole numbers.

Date

Beginning balance

Lease payment

Interest expense

Principal repayment

Closing balance

  1. Prepare the journal entries to record the inception of the lease on 1 July 2018.
  2. Prepare the journal entries on 30 June 2019.   
Answer #1
i) PV of the lease rent = (200000-25000)*3.6048 = $             6,30,840
PV of bargain purchse price = 300000*0.5674 = $             1,70,220
PV of the lease payments $             8,01,060
As the PV of lease payments equals the fair value of
machinery at a discount rate of 12%, 12% is the
implicit rate of the lease.
ii) Date Beginning balance Lease payment Interest expense Principal repayment Closing balance
06/30/2019 801060 175000 96127 78873 722187
06/30/2020 722187 175000 86662 88338 633850
06/30/2021 633850 175000 76062 98938 534912
06/30/2022 534912 175000 64189 110811 424101
06/30/2023 424101 474993 50892 424101 0
iii) Leased asset 801060
Lease liability 801060
iii) Lease liability 78873
Interest expense 96127
Insurance and maintenance expenses 25000
Cash 200000
Depreciation expense (630840/5) 126168
Accumulated depreciation (Leased asset) 126168